Find the vector form of the equation of the plane that passes through the point P0 = (1, −2, 3) (5)

and has normal vector ~n =< 3, 1, −1 >.


1
Expert's answer
2021-07-05T06:44:24-0400

P0(1, -2, 3)


normal vector n = 3i + j - k



Position vector of point P = i - 2j + 3k


The vector form of a plane that consists of a vector r0 and a normal vector n is given by



n . ( r - r0) = 0


So, the required equation of plane is



(3i + j - k ) . [r - ( i - 2j + 3k ) ] = 0